OpenStudy (anonymous):

What amount of heat is required to increase the temperature of 75.0 grams of gold from 150°C to 250°C

OpenStudy (anonymous):

Q=m*c*T, so change in temperature is 250-150=100, converting into Kelvins ==> 100+273=373, so our equation will be (Q is energy required to heat) Q=mass*c (specific heat of material)* T (change in temperature). We must consider whether it will change it's phase, if not which is likely then just use the formula given. C gold is 0.129 J/gm K. So finally, Q=75g*0.129J/gK*373=3608.775J

OpenStudy (anonymous):

I'm not sure about correctness of specific heat, so be careful, use what's given in your question. I think it should be divided by 100 further. As i suspect it's 0.129J over 100grams.